a. Will a 1200-watt hair dryer operate on a 120-volt line if the current is limited to 15 amperes by a safety fuse? Can two hairdryers operate on this line?


b. Choose from 3A, 5A, 13A, and 30A: Which fuse is the most suitable for a circuit in which a refrigerator rated at 2.0 kW is to be connected to a 220 V supply?


solution please thank you.


1
Expert's answer
2022-04-27T13:34:14-0400

(a)

"I=\frac{P}{V}=\frac{1200}{120}=10\:\rm A<15\: A"

One hair dryer will operate. Two -- no.

(b)

"I=\frac{P}{V}=\frac{2000}{220}=9.1\:\rm A"

The most suitable fuse for a circuit is 13 A.